OpenStudy (hope210):

will medal A garage door opener has a power rating of 350 watts. If the door is in operation for 30 seconds, how many joules of energy are used? (Hint: Remember that 1 watt = 1 joule/sec) a.350 J b.11.7 J C.10,500 J D.0.086 J

OpenStudy (hope210):

@logan13

OpenStudy (anonymous):

Power = Rate of work done P = W/t Where W = work done, t= time 350/30 = Work done

OpenStudy (radar):

Joules and "watt seconds" are equivalent. Simply multiply the power times the time.